OpenStudy (anonymous):

5x-y≤10 graph th inequality

OpenStudy (anonymous):

Hi You still there?

OpenStudy (anonymous):

First graph the line as if it is an equation 5x-y=10. Then work out which side makes it true. Do you need help to graph the line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What do you mean by work out what side makes it true

OpenStudy (anonymous):

Do you use the y=mx + c method?

OpenStudy (anonymous):

For drawing a line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so what does the equation look like if you rearrange it to make it y=mx+c? Can you do that?

OpenStudy (anonymous):

im truly not the brightest when it comes to math

OpenStudy (anonymous):

that's ok - I just needed to know how much help you need You want to rearrange 5x-y=10 so that you have y by itself on one side. first we want to get rid of the 5x, so what do we do with it?

OpenStudy (anonymous):

y=-5x+10

OpenStudy (anonymous):

no that isn't quite right

OpenStudy (anonymous):

5x-y=10 then subtract 5x off both sides -y=-5x+10

OpenStudy (anonymous):

You lost the minus sign

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

yes im just not sure what to do next

OpenStudy (anonymous):

ok so we now need to get rid of the - sign next to the y, so we multiply through by -1, which just means we change the sign on everything. that gives y=5x-10

OpenStudy (anonymous):

Do you need any of that explained?

OpenStudy (anonymous):

i get multipling by -1 and changing all the signs i still dont understand how im going to gra this

OpenStudy (anonymous):

Getting there. Now we want to know the slope of the line and the intercept so we can graph it. Can you tell me what the intercept is for this line?

OpenStudy (anonymous):

umm 10 not sure

OpenStudy (anonymous):

Not quite. It is -10. Remember the minus sign belongs to the number that comes after it. A way of thinking about the intercept is, we want to know what will y= when x=0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now what about the slope?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Yes great. Can you draw a line with a slope of 5 and an intercept of -10? I would find another point on the line myslef. Say x=4, what will y =?

OpenStudy (anonymous):

im not sure 8

OpenStudy (anonymous):

y=5x-10 y = 5 times 4 - 10 y = 20 - 10 y=10

OpenStudy (anonymous):

I'll do a sketch and we are nearly there. You should do one too and see if it looks like mine.

OpenStudy (anonymous):

but are we multiplying 5 by 4

OpenStudy (anonymous):

We did to find that one point to help us draw the graph

OpenStudy (anonymous):

An easier way to draw a graph, I think is to make a set of co-ordinate pairs that make the equation true, and then join them up. x y 0 -10 1 -5 2 0 3 5 4 10 5 15 Do you see what we are doing here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok - can you draw the line graph now? then we will deal with the <= sign

OpenStudy (anonymous):

I mean \[\le\]

OpenStudy (anonymous):

 Does it look like this?

OpenStudy (anonymous):